# UAV Path Planning for Object Observation with Quality Constraints: A Dynamic Programming Approach

### v1

@@ IV-C4 Simulation results
In Figure. 9, we compare the path length of our algorithms with that of the initial RS, NPF, GTSP methods and of the path which yields the maximum quality. The cases from RS, NPF, GTSP are all divided into two parts, based on the metrics that whether a path has already achieved the quality requirements. Then we compare the path length of these two parts separately after applying dynamic programming and calculate the reduced percentage of the path length.
The reduction magnitude also slows down distinctly as $q^{*}$ increases. When focusing on the plot of initial paths that are satisfied, generally the dynamic programming algorithm reduces RS paths to greater extent than NPF paths and GTSP paths by around $15\%$. The reduction in NPF paths is around $5\%$ and GTSP paths $3\%$. For those unsatisfied initial paths, our algorithm still functions in relatively smaller extent, regardless of the negative values when quality threshold rises to over 0.5. Since the original paths are deficient in observation quality, our algorithm can compensate for this deficiency at a relatively small cost of less than $10\%$. As for the contrast with the max quality path, the reduction rates climb further to around $24\%$ when $n$ is 5, $35\%$ for larger $n$. On these occasions, GTSP paths have larger adjustment space than RS and NPF paths.
@@
Given that the OCR approach is used to quantify the observation quality, the average word recognition accuracy is shown in Figure. 10 for each algorithm at different quality threshold. The straight and horizontal lines at the top of each subplot represent the accuracy of the maximum quality paths. This value, $\eta$, does not reach 1 because there probably exists random noise or other factors that may exert influence on photos taken and OCR processes. However, when comparing the average recognition accuracy of each of our algorithms with the maximum one, it can be discovered that the algorithms’ accuracy is extremely close to the product of maximum accuracy and quality thresholds. For example, when $q^{*}$ is 0.9, the algorithms’ outcomes are around 0.8 which is approximately $\eta*q^{*}$. From this perspective, the effectiveness of our algorithm is further validated.
Refer to caption (a) Refer to caption (b) Refer to caption (c)
Refer to caption (d) Refer to caption (e) Refer to caption (f)
Figure 9: Reduction of path length in Airsim compared with initial paths and max quality path by our algorithms
Refer to caption (a) Refer to caption (b) Refer to caption (c)
@@
Figure 10: Average recognition accuracy at different quality threshold and object number
IV-C5 Running time
@@
In addition to the approximation ratio and the effectiveness of the dynamic programming algorithm, the average processing time is presented in Table II. All of the data units are seconds. Generally, the time cost nearly doubles with the increase of $n$ after 10. When $n$ rises from 5 to 10, it reduces the discretization grid by half, and then the number of observation points of each object quadruples. As the data indicates, there is only minor difference in the running time among different visiting orders, owing to the fact that the running time is only related to the size of the dynamic programming table, which is unique since the problem complexity is the same.
@@
TABLE II: Running time for DP (in seconds),$\epsilon=0.5$
@@ IV-D Discussion
Our dynamic programming algorithm can both help the original paths to meet the quality requirement and reduce the length of the qualified ones. Provided there are more overlapping observation areas of objects, it can be anticipated a more significant reduction rate on the length. Apart from solving the above problem, our algorithm is also endowed with the capability to deal with the situation in which each object has a different weight and various observation quality demands. Suppose the weight of the objects to be ${w_{1},w_{2},...,w_{n}}$ and the quality constraint ${q^{*}_{1},q^{*}_{2},...,q^{*}_{n}}$. The generalization to different weight is achieved by multiplying $w_{i}$ to the observation quality at all observation points of $o_{i}$ before determining observing order and running the dynamic programming algorithm. We can eliminate the observation points of $o_{i}$ that do not yield the required quality $q^{*}_{i}$, leaving the rest observation points that all meet the condition. The elimination process can be adapted to different objects under $q^{*}_{i}$, hence the second generalization is achieved. Furthermore, in the presence of obstacles, the algorithm still works by adding the edges between the objects and observation points that are not blocked by the obstacles.
V Conclusion

### v2

@@ 5.3.4 Simulation results
The results of the first simulation experiments are given first. In Figure 10, we compare the path lengths of our algorithms with that of the baselines. The cases from RS, NPF, GTSP are all divided into two parts, based on the metrics that whether a path has already achieved the quality requirements. Figure 11 presents the percentage of path lengths that our algorithm can reduce compared to the baseline algorithm.
(a) path length, $n$ = 5
(b) path length, $n$ = 10
(c) path length, $n$ = 20
Figure 10: Path length of algorithms in the cylinder situation under different observation quality constraints
(a) path length, $n$ = 5
(b) path length, $n$ = 10
(c) path length, $n$ = 20
Figure 11: Reduction percentage of baseline path length in the cylinder situation
It is apparent that the MaxQ path remains the longest and is significantly longer than other paths, with differences ranging from 100 m to 300 m. Since the RS, NPF, and GTSP algorithms do not consider the observation quality, they are represented as horizontal lines in Figure 10. As $q^{*}$ increases, the path lengths of our algorithm gradually increase. When $n=5$, the path length increases from 410 m to 459 m, When $n=10$, it increases from approximately 490 m to 530 m. Correspondingly, the percentage of reduction in path length also gradually decreases. For baseline paths that already meet the observation quality requirements, our algorithm has a more pronounced effect in reducing the length of RS paths, about 15%, compared to reductions of approximately 6% for NPF paths and 3% for GTSP paths. In cases where the baseline path does not attain the observation quality, our algorithm still functions to a lesser extent, regardless of the negative values for larger $q^{*}$. This indicate that our algorithm compensates or the deficiencies in observation quality of the baseline paths at a modest cost of less than 10%. Compared to the MaxQ path, when $n=5$, the reduction percentage further increases to about 25%, and for larger $n$, it ranges from 30% to 40%.
@@
The average word recognition accuracy is shown in Figure 12 for each algorithm at different quality thresholds. The straight and horizontal lines at the top of each subplot represent the accuracy of the maximum quality paths (MaxQ). This value, $\eta$, does not reach 1 because there probably exists oscillation or other noise factors that may exert influence on photos taken and OCR processes. However, when comparing the average recognition accuracy of each of algorithms with the maximum one, the algorithms’ accuracy is extremely close to the product of maximum accuracy and the corresponding quality threshold. As an example, when $q^{*}$ is 0.9, the algorithms’ outcomes are around 0.8 which is approximately $\eta*q^{*}$.
(a) object number = 5
(b) object number = 10
(c) object number = 20
@@
Figure 12: Average recognition accuracy at different quality thresholds and object numbers
The simulation results in the virtual city environment are shown in Figure 13. The trends in path length and reduction ratios exhibit clear similarities with those discussed above. The reduction on the initial path length for RS is the highest, approximately 20%. In comparison, the reduction in NPF paths are around 7%, and about 5% on GTSP paths. Even with high-quality requirements of $q^{*}=0.9$, our algorithm increases the path length by only about 5% to meet the observation quality constraint (still achieving 10% on RS paths). Conversely, the reduction percentage on MaxQ path lengths dramatically increases to around 40%. Furthermore, Figure 13(c) shows that the average recognition accuracy is slightly higher than the results in the cylinder environment. The accuracy is approximately 65% when $q^{*}=0.3$, and gradually rises to about 85% when $q^{*}=0.9$. The reason is that objects in the 3D city environments are predominantly planar, thus the impact of deviating angles on perception quality is diminished, making it generally easier to achieve higher recognition accuracy.
Figure 13: Path lengths of algorithms, reduction percentage and recognition accuracy in the 3D city environment
5.3.5 Running time
@@
In addition to the approximation ratio and the effectiveness of the dynamic programming algorithm, the average processing time is presented in Table 2. All of the data units are seconds. Generally, the time cost nearly doubles with the increase of $n$ after 10. When $n$ rises from 5 to 10, it reduces the discretization grid by half, and then the number of observation points of each object quadruples. As the data indicates, there is only minor difference in the running time among different visiting orders, owing to the fact that the running time is only related to the size of the dynamic programming table, which is unique since the problem complexity is the same.
@@
Table 2: Running time for dynamic programming algorithm (in seconds), $\epsilon=0.5$
@@ 5.4 Discussion
Our dynamic programming algorithm can both help the original paths to meet the quality requirement and reduce the lengths of the qualified ones. Provided there are more overlapping observation areas of objects, it can be anticipated a more significant reduction rate on the lengths. In seeking a comparatively good observation quality, there is a worthwhile trade-off on the quality and the path length. The algorithm can seek the path that has $90\%$ observation quality but an evident reduction in the length cost. Based on these results, it can be drawn that GTSP paths yield the relatively best observation order because of the least improvement by the dynamic programming algorithm to achieve a certain quality threshold. Besides, GTSP paths also give the order that contribute to reduce the maximum quality path to the greatest extent, as compared with other heuristic paths.
Apart from solving the above problem, our algorithm is also endowed with the capability to deal with the situation in which each object has a different weight and various observation quality demands. Suppose the weight of the objects to be ${w_{1},w_{2},...,w_{n}}$ and the quality constraint ${q^{*}_{1},q^{*}_{2},...,q^{*}_{n}}$. The generalization to different weight is achieved by multiplying $w_{i}$ to the observation quality at all observation points of $o_{i}$ before determining observing order and running the dynamic programming algorithm. We can eliminate the observation points of $o_{i}$ that do not yield the required quality $q^{*}_{i}$, leaving the rest observation points that all meet the condition. The elimination process can be adapted to different objects under $q^{*}_{i}$, hence the second generalization is achieved. Furthermore, in the presence of obstacles, the algorithm still works by adding the edges between the objects and observation points that are not blocked by the obstacles.
6 Conclusion